OpenStudy (anonymous):

What is the common difference between successive terms in the sequence? 9, 2.5, –4, –10.5, –17, ... –11.5 –6.5 6.5 11.5

OpenStudy (anonymous):

(a term) - (term before) = (difference)

OpenStudy (anonymous):

So if we use the following notations, \(a_n\) - some Nth term. \(a_{n+1}\) - some term which is right after the Nth term. \(d\) - common differnce Then, the "difference" is by definition, \(a_{n+1}-a_n=d\)

OpenStudy (cal.lavender):

(a_n=a_1+(n-1)d a_2 implies 2.5=9+(1)d 2.5-9=d -6.5=d

OpenStudy (anonymous):

However, \(a_{n+1}-a_n=d\) Has to give you the same difference for all n, and if that is not the case, then your sequence is not an arithmetic sequence (if is a sequence at all), and no common difference is going to exist. (How can there be a common differnce if the difference varies?)